Profit is always calculated as revenue minus expenses. If a company’s expenses were $15 million, write a symbolic rule that expresses profit, P, (in millions of dollars) in terms of revenue, R, (in millions of dollars). Determine profit when revenue is $25 million. Determine Revenue that will produce a profit of $15 million dollars.

Answer by jim_thompson5910(35256)

Profit = Revenue - Expenses

P = R - E

P = R - 15

So the symbolic rule is P = R - 15

----------------------------------------------------------------

When revenue is $25 million, then

P = R - 15

P = 25 - 15

P = 10

profit is $10 million

-------------------------------------------------------

We want a profit of $15 million, so,

P = R - 15

15 = R - 15

15+15 = R

30 = R

R = 30

so we need a revenue of $30 million